### v2

# Orbital Paramagnetism without Density of States Enhancement in Nodal-line Semimetal ZrSiS

Abstract
Unconventional orbital paramagnetism without enhanced density of states was recently discovered in the nodal-line semimetal ZrSiS. We propose a novel interband mechanism, linked to the negative curvature of energy dispersions, which successfully accounts for the observed anomalous response. This negative curvature originates from energy variation along the nodal line, inherent in realistic nodal-line materials. Our results suggest that such orbital paramagnetism provides strong evidence for the presence of nodal lines in ZrSiS, and serves as a hallmark of other nodal-line materials.
Orbital magnetism is one of the fundamental properties of solids, rooted in the seminal work of Landau and Peierls on free-electron and tight-binding models [1, 2]. In Dirac electron systems such as bismuth and graphene, significant orbital diamagnetism arises from the interband effect of a magnetic field [3, 4, 5, 6, 7]. This discovery has shown that orbital magnetism is highly sensitive to the electronic structure, leading to extensive ongoing research into their relationship. Orbital magnetism is a bulk response that is also sensitive to the geometric properties encoded in the wave functions, such as nontrivial topology and quantum geometry. Weyl semimetals [3, 7, 8, 4, 5, 9, 9, 10, 11, 6, 12, 13], topological insulators [14, 15, 16, 17], and nodal-line semimetals [18, 19, 20, 21, 22, 23, 24, 7, 25] are prominent examples of these connections. Among these, nodal-line semimetals have been less characterized experimentally. Consequently, orbital magnetism is expected to offer a more direct, bulk-sensitive probe reflecting their characteristic electronic structures, complementing standard techniques such as angle-resolved photoemission spectroscopy [26, 27, 28], quantum oscillations [29, 30, 31], and transport studies [32, 33, 34, 35].
Orbital paramagnetism, as an unconventional orbital response, has recently been observed in the nodal-line semimetal ZrSiS at low temperatures, when a magnetic field is applied along the $C_{4}$ rotation axis [31]. This finding challenges the established understanding based on Larmor’s classical picture and Landau’s theory, both of which regard orbital magnetism as inherently diamagnetic. Furthermore, as the temperature increases, the paramagnetism decreases and eventually changes to diamagnetism at around 120 K, which is also an unexpected behavior. A similar paramagnetism has also been observed in another nodal-line semimetal, SrAs3 [36]. These observations in nodal-line semimetals suggest a possible connection between the presence of nodal lines and the emergence of such orbital paramagnetism, though its microscopic origin remains elusive.
At present, a few origins of orbital paramagnetism are known, including the Van Hove singularity [37, 38, 39], flatband [40, 41], and other mechanisms [42, 43]. In most of these cases, however, orbital paramagnetism is accompanied by a strong enhancement of density of states (DOS), and thus it is often masked by the large Pauli paramagnetism. In this Letter, we propose a new mechanism to explain the orbital paramagnetism without enhanced DOS. We first present a quantitative analysis of the orbital magnetic susceptibility in ZrSiS using the effective models based on the density functional theory (DFT) calculations, successfully elucidating the observed paramagnetism, temperature dependence, and anisotropic behaviors [31, 44]. To further understand the origin of this orbital paramagnetism, we derive a simple effective model. Our results show that the observed orbital paramagnetism without DOS enhancement is attributed to the interband effect associated with negative curvature in the energy dispersion. This negative curvature originates from the energy variation along the nodal line, which is inherent in realistic nodal-line materials. While dispersions with negative curvature can in principle occur in other systems, they emerge more naturally and systematically in nodal-line semimetals due to the presence of line-like band crossing. This novel mechanism suggests that such orbital paramagnetism provides strong evidence for the nodal lines not only in ZrSiS but also potentially in various other nodal-line materials.
DFT calculations and models.—
Figure 1: Positions of the nodal points in the Brillouin zone. The dimensionless wave numbers $\bar{k}_{x}=k_{x}a/(2\pi)$, $\bar{k}_{y}=k_{y}a/(2\pi)$, and $\bar{k}_{z}=k_{z}c/(2\pi)$ are shown. (a) Color indicates the energy of the middle of the gap $(E_{c}({\bm{k}})+E_{v}({\bm{k}}))/2-E_{F}$ with $E_{c}({\bm{k}})$ and $E_{v}({\bm{k}})$ being the energy of the bottom (top) of the conduction (valence) band at $\bm{k}$, and $E_{\rm F}$ is the Fermi energy. (b) Color indicates the magnitude of the gap due to the SOI, $E_{c}({\bm{k}})-E_{v}({\bm{k}})$. (c, d) Energy dispersions near $E_{\rm F}$ (c) along the nodal lines (i) and (ii), and (d) along the nodal line (iii). The valence (conduction) band is shown in solid (dashed) line.
ZrSiS has a $P4/nmm$ structure [45, 46] and has a set of nodal lines [47, 31, 48, 49]. To capture the characteristics of these nodal lines in detail, we perform DFT calculations using Quantum-ESPRESSO and Wannier90 packages [50, 51, 52, 53, 54, 55] with the lattice parameters of $a=3.55$ Å and $c=8.07$ Å [56]. The details of these calculations are shown in Supplemental Material (SM) [57]. Figure 1 shows the positions of the nodal points in the Brillouin zone, at which the bottom of the conduction band ($E_{c}({\bm{k}})$) and the top of the valence band ($E_{v}(\bm{k})$) are close to each other. The colors in Fig. 1(a) indicate the energy of the middle of the gap relative to the Fermi energy $E_{F}$, i.e., $(E_{c}(\bm{k})+E_{v}(\bm{k}))/2-E_{F}$, and those in Fig. 1 (b) indicate the magnitude of the gap $E_{c}(\bm{k})-E_{v}(\bm{k})$ due to the spin-orbit interaction (SOI). Four nodal lines (i) – (iv) are indicated, which are in the planes of $k_{z}=\pi/c$, $k_{z}=0$, $|k_{x}|=|k_{y}|$, and $k_{x}=0$ or $k_{y}=0$, respectively. Note that the nodal line protected by the nonsymmorphic symmetry [47] is omitted because it is approximately 1eV away from $E_{F}$. From Fig. 1(a), we can see that the nodal lines (i) and (ii) form closed lines in the Brillouin zone on the $k_{z}c=\pi$ and $k_{z}c=0$ planes, respectively, both of which cross $E_{F}$. As shown in Fig. 1 (c), the energy deviations around $E_{F}$ are approximately $\pm 0.25$ eV for (i) and $\pm 0.05$ eV for (ii), which play important roles in the later calculations. Figure 1(d) represents the energy of the valence and conduction bands along the nodal line (iii), where the Fermi level lies outside the Dirac mass gap. Nodal lines (iii) and (iv) form one-dimensional lines along the $k_{z}$ direction. While the nodal line (iii) is located close to $E_{F}$, (iv) is away from $E_{F}$, so that we neglect the nodal line (iv) in the following.
Considering these characteristics of ZrSiS, we introduce an effective model for the nodal lines (i) and (ii) as
${\mathcal{H}}^{\text{(X)}}=\biggl{[}\frac{\hbar^{2}}{2m^{*}}(k_{x}^{2}+k_{y}^{2})-\Delta\biggr{]}\sigma_{z}+\hbar v_{z}k_{z}\sigma_{x}+\eta^{\text{(X)}}(k_{x}^{2}-k_{y}^{2})\sigma_{0},$ (1)
where X=i or ii is the index for the nodal lines and $\sigma_{0}$, $\sigma_{x}$, and $\sigma_{z}$ are the identity and Pauli matrices. We have assumed a circular nodal line instead of the expected square nodal line for simplicity. This will be justified later by a comparison with the experimental data. The first term represents the two parabolic bands with masses of different signs. As shown in Fig. 2, the overlap of the bands is $2\Delta$. The second term in Eq. (1) hybridizes the two bands where $v_{z}$ represents the velocity in the $z$ direction. The $\bm{k}$} points on the nodal line satisfy $\frac{\hbar^{2}}{2m^{*}}(k_{x}^{2}+k_{y}^{2})-\Delta=0$ and $k_{z}=0$, indicating that the nodal line forms a circle in the $k_{x}$-$k_{y}$ planes at $E_{F}=0$. Finite $k_{z}$ opens gaps along the nodal line. The last term introduces variation of energy along the nodal line, and the energy deviation is $\pm 2\eta^{\text{(X)}}k_{\rm R}^{2}$ with $k_{\rm R}=\sqrt{2m^{*}\Delta}/\hbar$ being the radius of the nodal line. We will see below that this variation play a crucial role for the emergence of orbital paramagnetism. Similar energy variation has been discussed in Refs. [42, 58, 59, 60, 25]. In the following, we omit the superscript of $\eta$ when a distinction is not necessary. The DOS is also shown in Fig. 2 for $\tilde{\eta}=0$ and $0.2$, where $\tilde{\eta}$ is a dimensionless parameter $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta$. The energy variation results in the emergence of DOS around $E=E_{F}$. This variation also gives negative Gauss curvature of the $E$-$k_{x}$-$k_{y}$ surface at $(k_{x},k_{y},k_{z})=(0,\pm k_{\rm R},k_{z0})$ for the valence band and at $(\pm k_{\rm R},0,k_{z0})$ for the conduction band for $k_{z0}\neq 0$, i.e., in the presence of gaps. The curvature is illustrated in detail in SM [57].
[…]
Note that the effective model of Eq. (1) does not include the SOI effect. As discussed later, SOI hardly affects our results.
The nodal line (iii) is approximately regarded as 2D Dirac electrons, governed by
$\mathcal{H}^{\text{(iii)}}=\hbar v_{\rm Dirac}(k_{x}\sigma_{x}+k_{y}\sigma_{y})+\Delta_{\rm SOI}\sigma_{z}+\varepsilon_{0}(k_{z}),$ (2)
where $\varepsilon_{0}(k_{z})$ is the energy shift and $\Delta_{\rm SOI}$ represents the gap originating from SOI. We discuss the orbital magnetic susceptibility using models (1) and (2) in the following.
Orbital magnetic susceptibility.— The orbital magnetic susceptibility in a magnetic field in the $z$ direction ($\chi_{z}$) is generally given by [61]
$\chi_{z}=\frac{e^{2}}{\hbar^{2}}k_{B}T\sum_{n{\bm{k}}}{\rm Tr}\,\gamma_{x}{\mathcal{G}}\gamma_{y}{\mathcal{G}}\gamma_{x}{\mathcal{G}}\gamma_{y}\mathcal{G},$ (3)
where ${\mathcal{G}}=\mathcal{G}({\bm{k}},i\varepsilon_{n})=[i\varepsilon_{n}-\mathcal{H}^{\text{(X)}}+\mu]^{-1}$ is the thermal Green’s function, $\varepsilon_{n}=(2n+1)\pi k_{\rm B}T$ ($n\in\mathbb{Z})$ is the Matsubara frequency, and $\gamma_{i}$ is the velocity operator of in the $i$ ($i=x,y,z$) direction defined by $\gamma_{i}=\frac{\partial{\mathcal{H}}^{\text{(X)}}}{\partial k_{i}}$. Equation (3) is a general formula for noninteracting systems and incorporates the spin degeneracy. The orbital magnetic susceptibilities in other directions, $\chi_{x}$ and $\chi_{y}$, are obtained by a cyclic replacement of $x$, $y$, and $z$. We numerically evaluate Eq. (3) for the model Eq. (1), employing the quasi-Monte Carlo method [62, 63, 64, 65, 66] for $\bm{k}$} summation, and the sparse-ir method [67, 68, 69] for Matsubara summation. The momentum cutoffs are $\Lambda_{R}=1000k_{R}$ in the radial direction and $\Lambda_{z}=(l_{z})^{-1}$ in the $k_{z}$ direction with $l_{z}=\hbar v_{z}/\Delta$. The Matsubara summation includes a sufficient number of frequencies corresponding to the energy cutoff $\varepsilon_{\rm cut}=100\Delta$  [^1] In our approach, we adopt a finite but sufficiently large cutoff for Matsubara summation. Correspondingly, we choose the $\bm{k}$} integral cutoff to be much larger than the Matsubara frequency cutoff. This method confirms that the numerical integration can be carried out efficiently and with sufficient accuracy, when compared with the analytical results obtained in Ref. [24] and SM [57].
The results for $\chi_{z}$ and $\chi_{x}$ are shown in Fig. 3 (a) and (b), respectively, for several values of $\tilde{\eta}$. For the case without the energy variation ($\tilde{\eta}=0$), $\chi_{z}$ is diamagnetic for every $\mu$ while $\chi_{x}$ has a sharp peak at $\mu=0$ originating from the interband effect of the 2D Dirac electrons [6]. We confirmed that our calculation reproduces the previous results [24] in the limit of $\Lambda_{\rm R}\to\infty$ and $\Lambda_{z}\to\infty$ [see SM[57] for detail].
Finite $\eta$ gives a significant effect on $\chi_{z}$: $\chi_{z}$ has a broad peak around $\mu=0$, whose width is approximately $2\eta k_{\rm R}^{2}$. Furthermore, for $\tilde{\eta}>0.1$, the value at the peak is positive, meaning orbital paramagnetism. The inset of Fig. 3 (a) shows the Landau-Peierls (LP) contribution [1, 2, 38, 71], or the intraband contribution, which is negative for all values of $\mu$. Therefore, we conclude that the obtained orbital paramagnetism near $\mu=0$ is due to an interband effect. As we noted before, while the orbital paramagnetism is usually accompanied by large DOS, the present result suggests an interband orbital paramagnetism without the enhancement of the DOS. The presence of two energetically close bands involving negative curvature is a crucial factor in inducing orbital paramagnetism in this system. While such a configuration can, in principle, arise in other materials, it remains primarily associated with nodal-line semimetals and is seldom realized elsewhere. The mechanism for this paramagnetism is discussed in detail later.
Figure 3: (a) Orbital magnetic susceptibility $\chi_{z}$ as a function of chemical potential when $k_{B}T=0.01\Delta$ for several values of $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta$. (b) Orbital magnetic susceptibility $\chi_{x}$ when $k_{B}T=0.05\Delta$. They are normalized by $\chi_{0}\equiv\frac{e^{2}}{\hbar^{2}}Vl_{z}\Delta$ with $V$ being the volume and $l_{z}=\hbar v_{z}/\Delta$. Inset: LP contribution for $\tilde{\eta}=0.2$ and $k_{B}T=0.1\Delta$.
As shown in Fig. 3 (b), $\chi_{x}$ is not an even function of the chemical potential for finite $\eta$. We find the relations $\chi_{x}(\mu,\eta)=\chi_{x}(-\mu,-\eta)$ and $\chi_{x}(\mu,\eta)=\chi_{y}(-\mu,\eta)$ from the analytical expressions, shown in SM [57]. This behavior is understood as follows. At each point on the nodal line, a Dirac dispersion $\propto\pm\sqrt{q_{x}^{2}+q_{y}^{2}}$ is formed, where $q_{x}$-$q_{y}$ plane is perpendicular to the tangent of the nodal line at the point. Therefore, when the magnetic field is parallel to the tangent of the nodal line, the delta function–like orbital diamagnetism ($\propto-\delta(\mu-\varepsilon_{0}))$ [61, 6] is negatively maximized with $\varepsilon_{0}$ being the energy of the Dirac point. As we can see from the symbol $\times$ in Fig. 2, when $\eta$ is positive, the energy of the nodal point at which the tangent of the nodal line is parallel to $\tilde{k}_{x}$ is negative, i.e., $\varepsilon_{0}<0$. Thus, $\chi_{x}$ takes the negatively maximal value at $\mu=\varepsilon_{0}<0$. The magnitude of $\chi_{x}$ becomes smaller as $\eta$ increases, since the region of the nodal line parallel to $\tilde{k}_{x}$ becomes smaller.
Comparison with experiments.— We evaluate the orbital magnetic susceptibility for ZrSiS using the parameters obtained by the DFT calculations. For $\chi_{z}$, we consider $\chi_{z}^{\rm tot}=\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}+\chi_{z}^{\text{(iii)}}+\chi_{\rm Pauli}$, where $\chi_{z}^{\text{(X)}}$ represents the contribution from the nodal line X (X=i, ii, iii), and the Pauli paramagnetism $\chi_{\rm Pauli}=0.125\times 10^{-4}$ emu/mol [31] is included. We evaluate $\chi_{z}^{\text{(i)}}$ and $\chi_{z}^{\text{(ii)}}$ from Eq. (1). $\chi_{z}^{\text{(iii)}}$ is obtained from the model Eq. (2) by the $k_{z}$ average of the known result for the 2D Dirac electron system [38]
$\chi_{2D}(k_{z})=\frac{e^{2}v_{\rm Dirac}^{2}}{6\pi}\frac{f(\Delta_{\rm SOI},-\varepsilon_{0}(k_{z}))-f(-\Delta_{\rm SOI},-\varepsilon_{0}(k_{z}))}{\Delta_{\rm SOI}}$ (4)
as $\chi_{z}^{\text{(iii)}}=\frac{V}{2\pi}\int dk_{z}\chi_{2D}(k_{z})$ where $f(\varepsilon,\mu)$ is the Fermi distribution function.
From the DFT calculations [see Fig. 1], we obtain $\tilde{\eta}^{\text{(i)}}=0.31$ for (i), $\tilde{\eta}^{\text{(ii)}}=0.06$ for (ii), and $\Delta_{\text{SOI}}=15$ meV and $\varepsilon_{0}(k_{z})=(-0.025-0.07|k_{z}|c/\pi)$ eV for (iii), respectively. Note that these parameters ensure that the Fermi level lies outside the Dirac mass gap along (iii). The other parameters are adjusted to fit the experimental data [31], which are $\Delta=0.8$ eV, $m^{*}/m_{0}=0.12$, $v_{z}/c_{0}=4\times 10^{-4}$, and $v_{\rm Dirac}/c_{0}=9\times 10^{-4}$ with $m_{0}$ and $c_{0}$ being the bare electron mass and the speed of light in vacuum, respectively. Figure 4 shows the results of $\chi_{z}^{\text{tot}}$ and $\chi_{x}^{\text{tot}}$, where $\chi_{z}^{\rm tot}$ agrees well with the experimental data quantitatively and $\chi_{x}^{\rm tot}$ captures the observed trends qualitatively.
We now discuss each contribution in $\chi_{z}^{\text{tot}}$. Since we have $\tilde{\eta}^{\text{(i)}}=0.31$ and $\tilde{\eta}^{\text{(ii)}}=0.06$, near $T=0$, $\chi_{z}^{\text{(i)}}$ is positive, while $\chi_{z}^{\text{(ii)}}$ is small and negative, as shown in Fig. 3(a). Their sum is $\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}=0.26\times 10^{-4}$ emu/mol, which is the main contribution for orbital paramagnetism near $T=0$. For temperatures below 300 K, $\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}$ is almost constant. The temperature dependence of $\chi_{z}^{\text{tot}}$ is attributed to $\chi_{z}^{\text{(iii)}}$. At low temperatures, $\chi_{z}^{\text{(iii)}}$ is negative due to Dirac electrons in the $k_{x}$-$k_{y}$ plane but small because the chemical potential is slightly outside the gap [see Fig. 1 (d)]. As the temperature increases, the diamagnetism from $\chi_{z}^{\text{(iii)}}$ grows because of the smearing as expressed by Eq. (4). This leads to negative $\chi_{z}^{\text{tot}}$ when $T>120$ K.
Note that the effect of the gap due to SOI is neglected in the model Eq. (1) used in the above calculations. This treatment is justified, as the gap is small ($\lesssim 0.03$ eV) compared with $\Delta$ (=0.8=0.8 eV), which limits its effect on $\chi_{z}$ in Fig. 3 to the narrow range $|\mu|/\Delta\lesssim 0.04$. Explicit evaluation of the effect of gap opening is provided in SM [57]. Even in the presence of band hybridization, the residual nodal-line features continue to play a crucial role, leaving a strong imprint on the interband response and still leading to paramagnetism.
Next, we consider $\chi_{x}^{\text{tot}}=\chi_{x}^{\text{(i)}}+\chi_{x}^{\text{(ii)}}+\chi_{\rm Pauli}$, where we have assumed $\chi_{x}^{\text{(iii)}}=0$ since the magnetic field is perpendicular to the Dirac cone along the nodal line (iii). We can see that the negative value of $\chi_{x}^{\text{tot}}$ in Fig. 4 originates from $\chi_{x}$ of the nodal lines (i) and (ii) shown in Fig. 3 (b). Furthermore, we find that $\chi_{x}^{\text{(i)}}+\chi_{x}^{\text{(ii)}}$ exhibits a much weaker temperature dependence than $\chi_{z}^{\text{(iii)}}$ discussed above, although in both cases the nodal lines include segments parallel to the magnetic field. The essential difference between the nodal lines (i) and (ii), on the one hand, and (iii), on the other hand, is the energy of the Dirac points relative to $E_{F}$: (i) and (ii) cross $E_{F}$, whereas (iii) lies below $E_{F}$ [See Fig. 1 (c) and (d)]. This difference leads to the difference of the temperature dependencies between $\chi_{x}^{\rm(i)}+\chi_{x}^{\rm(ii)}$ and $\chi_{z}^{\rm(iii)}$.
Figure 4: Temperature dependence of $\chi_{z}^{\rm tot}$ and $\chi_{x}^{\text{\rm tot}}$. Experimental data for $\chi_{z}^{\rm exp}(\bigcirc)$ and $\chi_{x}^{\rm exp}(\square)$ [31] are also shown. The enhancement of magnetic susceptibility $\chi_{z}$ below 10 K is owing to localized magnetic impurities [31], not the orbital effect.
The models (1) and (2), based on DFT calculations, capture the essential features of the material. Our calculations achieve quantitative agreement with the experimentally measured $\chi_{z}$ and successfully reproduce the characteristic behavior of $\chi_{x}$. These results support the validity of the circular approximation for the nodal-line geometry and confirm that the energy variation and negative curvature in the dispersion are key ingredients for the observed orbital paramagnetism. We note that this quantitative agreement can be attributed to the weak dependence of $\chi_{z}^{\rm(i)}$ and $\chi_{z}^{\rm(ii)}$ on the chemical potential $\mu$, as shown in Fig. 3(a), which leads to robustness against small variations in the omitted details. In contrast, as illustrated in Fig. 3(b), $\chi_{x}^{\rm(i)}$ and $\chi_{x}^{\rm(ii)}$ exhibit stronger $\mu$ dependence, suggesting that more realistic models for nodal line (i) and (ii) would be necessary to reproduce the experimental results with higher precision.
Effective model and discussions.— To understand the mechanism underlying the interband orbital paramagnetism, which is essential for $\chi_{z}$ in ZrSiS, we introduce an effective model,
$H=\alpha\frac{k_{x}^{2}-k_{y}^{2}}{2}\sigma_{0}+\alpha\frac{k_{x}^{2}+k_{y}^{2}}{2}\sigma_{z}+d\sigma_{x},$ (5)
where $\alpha$ and $d$ represent the energy scale and band hybridization, respectively. This model is obtained from a small $(k_{x},k_{y})$ expansion of Eq. (1) around $(k_{x},k_{y},k_{z})=(\pm k_{\rm R},0,k_{z0})$, for some fixed $k_{z0}\neq 0$. Figure 5 shows the orbital magnetic susceptibility in the $z$ direction, $\chi_{\rm orb}$, calculated by Eq. (3). The LP (intraband) contribution $\chi_{\rm LP}$ [1, 2, 38, 71] and the interband contribution ($\chi_{\rm inter}=\chi_{\rm orb}-\chi_{\rm LP}$) are also shown in Fig. 5. Our results show that this model exhibits orbital paramagnetism near $\mu/d\sim 0$ where the ground state is insulating, which is not explained by the intraband effect. The dependence on the chemical potential resembles that observed in Fig. 3(a), clearly showing that the orbital paramagnetism in Fig. 3(a) originates from the interband effect between the saddle points. This new mechanism stands in stark contrast to orbital paramagnetism arising from the Van Hove singularity in single-band 2D systems [37, 38] or flatband systems [40, 41], as it does not rely on divergent DOS or even the presence of Fermi surfaces. The differences between these mechanisms of orbital paramagnetism are summarized in Table 1.
Table 1: Comparison of orbital paramagnetism mechanisms across different systems
System Type Fermi surface DOS
Van Hove
singularity [37]
Intraband Necessary Divergent
Flatband [41] Interband Necessary Divergent
Nodal line (present) Interband Not required Non-divergent
Figure 5: Total orbital magnetic susceptibility $\chi_{\rm orb}$, LP contribution $\chi_{\rm LP}$, and the interband contribution $\chi_{\rm inter}\equiv\chi_{\rm orb}-\chi_{\rm LP}$ of the saddle point model Eq. (5) in units of $\chi_{1}=e^{2}/\hbar^{2}$ with $k_{\rm B}T=1/40$, where we have used the unit of $\alpha=d=1$. Numerical integration has been conducted on a region $|k_{x}|\leq k_{c}$ and $|k_{y}|\leq k_{c}$ with $k_{c}=4$, showing weak cutoff dependence. Inset: Energy dispersions for Eq. (5).
Summary.— We have studied the orbital magnetism in ZrSiS based on the DFT calculations and an effective model. Our results elucidate three anomalies observed in the orbital magnetism, the large orbital paramagnetism without DOS enhancement, temperature dependence, and paramagnetic-to-diamagnetic anisotropy. We have found that the orbital paramagnetism in the $C_{4}$ axial direction ($k_{z}$ direction) at low temperatures is attributed to an interband effect. This mechanism is novel in that the orbital paramagnetism is not accompanied by an enhancement of DOS. Therefore, the Pauli paramagnetism is suppressed, enabling the clear observation of orbital paramagnetism. This interband effect is understood in terms of a simple effective two-band model. In this model, the negative curvature, which arises from variation of the energy of the node points along the nodal line in the material-specific model, induces the anomalous orbital response. This variation, which persists even when a small band hybridization is present, is typical of nodal-line semimetals and difficult to achieve in conventional systems. Orbital paramagnetism is one of the most prominent features of nodal-line semimetals, and importantly, it is experimentally accessible.
Acknowledgements.
We are grateful to Y. Suzumura, M. Hayashi, Z. Hiroi, T. Osada, N. Tsuji, N. Kawashima, and H. Shinaoka for their insightful discussions. This work was supported by Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science (No. JP22J15355, No. JP18H01162, No. JP21H05191, No. JP18K03482, No. JP17H02912, No. JP21H01003, No. JP22K03447, and No. JP23H04869).